D Question 11 1 pts A study reports the following final notation: F(2, 37) = 9.00,p<.05; how many total participants were inv I need help with questions 11 & 12 please.
Answer #1

Question 11

We are given

df1 = 2

df2 = 37

Total df = df1 + df2 = 2 + 37 = 39

n = Total df + 1 = 39 + 1 = 40

n = 40

Answer: 40

Question 12

We have

F = MS between / MS within

We are given F = 9.00, MS within = 5,

So we have

9 = MS between / 5

MS between = 9*5 = 45

Answer: 45